# ReUseIt: Synthesizing Reusable AI Agent Workflows for Web Automation

### v1

@@ 4.2. Agent Execution Guided by Synthesized Workflow (DG2)
Figure 6. Agent execution guided by the synthesized workflow. At actions where pre-/post-condition checks exist in the workflow, ReUseIt captures the webpage screenshot and does a Q&A to determine if the condition check is met before/after this action. If the answer is ‘Yes’, the agent proceeds to the next action, but if not, it retries executing this action according to the fallback actions. The agent moves on the the next action if the condition check is met after retries. However, if the retries fail, ReUseIt notifies the user by providing a challenge analysis to help them understand the challenge and provide guidance. User guidance is used to iteratively update the condition checks and fallback actions in the workflow.
Condition check and agent self-recovery. With synthesized workflows, ReUseIt performs condition checks before/after each agent action. For each check, it uses the condition check prompt to evaluate if the expected conditions are met. First, ReUseIt captures the webpage screenshot before/after each action. It then pairs the screenshot with the condition check, and asks the LLM to produce a “Yes/No” answer along with an explanation. If the condition is met, the agent moves on to the next action; if not, it self-recovers by re-executing the failed step up to three times according to the fallback actions.
User notification. If the agent’s self-recovery fails, ReUseIt identifies and explains the issue to notify the user. It uses the challenge identification and explanation prompt to instruct the LLM to analyze the condition check, as well as agent messages and screenshots from all attempts, in order to produce three aspects of content: (1) where the failure occurred; (2) why the agent failed; (3) what agent behaviors caused the failure. The LLM first refers to the condition check to locate at which step the agent failed. Next, it explains why the agent’s actions were considered failed based on the conditions that were not met. In addition, it summarizes agent messages and webpage screenshots and analyzes what specific actions led to the failure. Lastly, ReUseIt offers actionable tips to help the user provide guidance to unblock the agent’s progress. It assembles the actionable guidance prompt to guide the LLM to summarize the fallback actions at failed steps and outline insights into how to help the agent recover that the user can reference.
@@ 4.3. Implementation
Web agent framework. Our system is built on the Magentic-UI web agent framework (Mozannar et al., 2025d). We use Magentic-UI’s web surfer (Mozannar et al., 2025b) for web browsing, interaction, and navigation, as well as the same tool definitions (e.g., visit URL, web search, scroll down/up, click, type) (Mozannar et al., 2025c) to define the web surfer’s functionality.
LLMs. ReUseIt has been tested with OpenAI GPT-4o for webpage parsing and action reasoning of the web surfer. The reasoned actions are performed on the webpage using Playwright (Microsoft, 2025a). The LLMs for workflow synthesis, condition check on webpage states, and user notification were tested with GPT-4o.
Workflow synthesis time. The execution time for each task depends on the number of steps. We report the mean and standard deviation of per-task execution time for the formative study tasks in Table 2. To synthesize the ReUseIt Workflow, each original and variation task is executed five times, so the workflow synthesis time is about the total number of executions multiplied by the per-task execution time.
Table 2. Execution and workflow synthesis time of user study tasks (mins:secs).
@@
1 Search the ¡flight type¿ ¡cabin type¿ ticket from ¡departure city¿ to ¡destination city¿ on ¡date¿ for ¡number of passengers¿ on ¡website¿ $2{:}38\pm 0{:}41$ 52:4052{:}40
2 Go to ¡website¿ to find ¡home filtering¿ ¡home type¿ for ¡rent or buy¿ in ¡city¿. Scrape the information from top ¡num listings¿ $1{:}20\pm 0{:}47$ 26:4026{:}40
3 Go to ¡website¿ to find the first ¡num pets¿ ¡pet filtering¿ ¡pet type¿’s ¡pet information¿ that can be adopted near ¡city¿ $1{:}33\pm 0{:}31$ 31:0031{:}00
4 Search the lowest price of ¡product¿ with ¡attributes¿ by applying ¡discount¿ on ¡website¿ $1{:}11\pm 0{:}38$ 23:4023{:}40
5 Search the ¡paper title¿ on Google Scholar. Find the papers authored by ¡author of interest¿ of the searched article. Retrieve the most recent paper according to ¡sorting criteria¿ and summarize its abstract $0{:}46\pm 0{:}20$ 15:2015{:}20
6 Search ¡news topic¿ from ¡website¿, and summarize its author’s recent articles $1{:}09\pm 0{:}39$ 23:0023{:}00
@@ 5.1. Evaluation Setup
We randomly sampled fifteen tasks from the Skyvern web benchmark (Singh, 2025) for the benchmark evaluation (task details in Table 3). For each task, we automatically generated variation tasks (using the method introduced in Section 4.1), executed the original and variation tasks (a task family) five times, and then synthesized a workflow for each task family. We compared the agent guided by Task + our synthesized ReUseIt Workflow (e.g., Figure 7) against three baselines: Task-Only, Task + Success-Traces, and Task + Magentic-UI Plan. We used the average success rate ($SR$) as the evaluation metric (as introduced in Section 3.2). Following WebVoyager (He et al., 2024), we employ LLM-as-a-judge to assess execution success. A GPT-4o evaluator takes the task description, the final three screenshots prior to task termination, and the agent’s answer, and returns a binary success decision accompanied by a rationale. To validate the performance of this LLM judge, we randomly audited 45 judgments (15%) with a human rater. The LLM-judge achieved $\textrm{Acc}=0.778$, $\textrm{Precision}=0.852$, $\textrm{Recall}=0.793$, $\textrm{F1}=0.821$. Agreement between the LLM judge and human rater was moderate-substantial with $\textrm{Cohen's }\kappa=0.528$.
@@ 5.2. Evaluation Results
Figure 8 presents the benchmark evaluation results for all fifteen task families. We observe that our synthesized workflow has improved the $SR$ by a large margin compared to the baseline methods. Averaging across tasks, the $SR$s are: Task + ReUseIt Workflow ($70.1\%\pm 16.4\%$), Task + Magentic-UI Plan ($48.6\%\pm 12.9\%$), Task + Success-Traces ($41.4\%\pm 14.8\%$), Task-Only ($24.2\%\pm 13.2\%$). ReUseIt Workflow improves $SR$s by about $21.5\%$ over Task + Magentic-UI Plan, $28.7\%$ over Task + Success-Traces, and $45.9\%$ over Task-Only.
From the results, we find that providing the agent with successful traces has been effective in improving the $SR$s on the original task, but less effective on the variation tasks compared to Magentic-UI Plan. The largest improvements come from adding execution guards in the ReUseIt Workflow. The condition checks and fallback actions have significantly increased agent reusability by guiding the agent to reflect on actions and self-recover from failures. These benefits also improved reliability in executing the same and similar tasks. Additionally, our evaluation spans multiple variation tasks, so higher success rates indicate better generalizability to handle value adjustments and webpage UI differences.
Figure 8. Benchmark evaluation results on fifteen web automation tasks and their variations (avg $\pm$ std success rate).
@@ 6. User Study
To understand how our system affects the user experience when interacting with web agents, we conducted a user study with the following research questions:
@@ 6.1. Study Design
Procedure. The onboarding session ($\sim$10 mins), involving consent, introduction, and an example task walk-through, followed the same procedure as the formative study. After it, each participant examined two task families covering six agent executions ($\sim$30 mins). We shuffled the task order and ensured each task family was examined by three different participants. In each task, they were provided with a slide deck containing the task description, the screen recording of agent actions, the user notification message, and the workflow showing a sequence of actions the agent followed. After examining all this information, participants were asked to verbally analyze why and how the agent failed or successfully completed the tasks, evaluate the helpfulness of the user notification message to understand the agent’s issues, and (optionally) offer guidance by either writing a short instruction or typing their edits in the workflow. We recorded the desktop screen and transcribed the audio. After the task, participants completed a post-study questionnaire and had a semi-structured interview ($\sim$20 mins). The interview questions can be found in Appendix B.1. Each participant was paid a \$30 gift card to acknowledge their participation.
@@ 6.2. Quantitative Measures
Follow-Up Experiment (S2-RQ1). Figure 9 shows the follow-up experiment results. In tasks “Product Price Tracking” and “Publication Search,” participants were satisfied with the ReUseIt Workflow and did not improve it, and we plotted the success rate averaged across only the users who offered guidance (Task + ReUseIt Workflow User). See Figure 7 for an example workflow updated with user guidance (blue text). Appendix B.2 reports the full results for each user. Across tasks, the success rates for Task + ReUseIt Workflow and Task + ReUseIt Workflow User were $86.5\%\pm 9.9\%$ and $88.6\%\pm 6.1\%$. These results show that (1) ReUseIt Workflow can significantly improve the agent’s task execution reliability compared to the baseline methods and (2) user guidance can further help improve the workflow for better agent execution performance on repetitive tasks.
Figure 9. User study follow-up experiment results on six web automation tasks and their variations (avg $\pm$ std success rate).
@@
Accompanying the question on required user guidance, we conducted an analysis of user guidance statistics to compare the number and length of user guidance. We received 9, 17, and 17 pieces of user guidance to ReUseIt Workflow, Magentic-UI Plan, and Task-Only, respectively. The average words ($\pm$ std) in user guidance for the three conditions were 56.11 $\pm$ 44.59, 43.24 $\pm$ 48.18, and 43.41 $\pm$ 50.02. These results indicate that the execution guards in the ReUseIt Workflow have effectively improved the interpretability of agent actions/issues and reduced the guidance burden on users.
Figure 10. User study post-study questionnaire results.
@@ 6.3. Qualitative Findings
From the analysis of interview data, we obtain the following key findings centered around S2-RQ2 to S2-RQ4. Overall, we found that the ReUseIt ’s synthesized workflows have helped users to spot and fix agent issues and augmented their willingness to adopt web agents.
Condition checks helped users to identify issues, issue analysis helped them to understand challenges (S2-RQ2). Participants emphasized that ReUseIt Workflow’s condition checks helped them anticipate and localize failures without exhaustive monitoring. S2-P3 noted, “If you know where the model’s gonna fail from the condition checks [in ReUseIt Workflow], you don’t have to watch the agent very carefully and examine everything. You kind of know when the error’s gonna be. Otherwise, there’s always some possibility of making an error at any step. So definitely helps you pinpoint where things are gonna go wrong.” They also commented that the user notification guided by ReUseIt Workflow clarified the reason why the agent failed, which helped to make issue diagnosis easier. For instance, S2-P4 said that “With ReUseIt Workflow, I would understand a little bit better why,” and S2-P3 emphasized the value of this detail: “For ReUseIt Workflow, I feel like even if it gets it wrong, it gives you so much detail that you can exactly intervene on the spot that you want to.”
@@
User notification with challenge identification and guidance tips allowed precise and low-effort user guidance (S2-RQ3). Participants described the user notification with ReUseIt Workflow broadly covered the agent’s challenges that supported targeted edits. As S2-P1 explained that “The messages are like a super set of the challenges faced by the agent. I always find that my instruction is a subset of the suggested ideas. And also, I think it’s important that I match the suggested ideas with the actual workflow of the agent because some of the suggested content is already covered by the workflow, so I don’t have to repeat it. Instead, I need to think about why the existing workflow is not good enough.” In addition, S2-P2 underscored the usefulness of actionable guidance, saying that “I do think that the actionable guidance says something about potentially what are some of the next steps, and after reviewing the workflow myself and then editing the prompts, I think the actionable guidance does have a good contribution.” The ReUseIt Workflow, in turn, supported user understanding of agent challenges for them to provide more effective guidance (S2-P1, S2-P8). For example, S2-P8 said that “These [condition checks and fallback actions] actually give me not only hints, but also much guidance on how I should train an agent, or what my approach should be.”
@@
Reliable and controllable web agents improved user willingness for practical use (S2-RQ4). Participants highlighted that the ReUseIt ’s workflows made reuse feasible because the agent behaved more consistently and could be steered through explicit condition checks. As S2-P2 put it, “If I wanna reuse it [the agent] reliably, then I would consider being very extensive with my workflow and being very consistent about it… so that I know how it’s supposed to work.” S2-P4 further commented on the connection between user trust and agent reliability, saying, “I think I have a better chance of getting better results just because it [the agent] has… ways of approaching the task, so I think I would trust it more. Especially if I had a certain level of control in terms of the conditions and the fallbacks, I would use it even more just to make sure that the result that I’m getting is the proper one.”

### v2

@@ 4.2. Agent Execution Guided by Synthesized Workflow (DG2)
Figure 6. Agent execution guided by the synthesized workflow. At actions where pre-/post-condition checks exist in the workflow, ReUseIt captures the webpage screenshot and does a Q&A to determine if the condition check is met before/after this action. If the answer is “Yes”, the agent proceeds to the next action, but if not, it retries executing this action according to the fallback actions. The agent moves on to the next action if the condition check is met after retries. However, if the retries fail, ReUseIt notifies the user by providing a challenge analysis to help them understand the issue and provide guidance. User guidance is used to iteratively update the condition checks and fallback actions in the workflow.
This figure shows a flowchart for agent execution guided by the synthesized workflow. In the left column, ReUseIt performs ”Condition Check” before/after step i by comparing the webpage screenshot and the synthesized condition checks to produce a Yes/No answer. If Yes, the agent moves on to step i+1; if No, the agent is guided to perform ”Agent Self-Recovery”, which is shown in the middle column. To self-recover, the agent retries step i according to the synthesized fallback actions and does the condition check again after the retries. If the condition is met, the agent continues to step i+1, but if not, the agent needs to send a ”User Notification” message, as shown in the right column. In this message, the agent identifies and explains its challenge, as well as offers the user actionable guidance that can help the user offer feedback to unblock the challenge. The user feedback is used to iteratively update the synthesized condition checks and fallback actions.
@@
Condition check and agent self-recovery according to fallback actions. With synthesized workflows, ReUseIt performs condition checks before/after each agent action. For each check, it uses the condition check prompt to evaluate if the expected conditions are met. First, ReUseIt captures the webpage screenshot before/after each action. It then pairs the screenshot with the condition check, and asks the LLM to produce a “Yes/No” answer along with an explanation. If the condition is met, the agent moves on to the next action; if not, it self-recovers by re-executing the failed step up to three times according to the fallback actions.
User notification. If the agent’s self-recovery fails, ReUseIt identifies and explains the issue to notify the user. It uses the challenge identification and explanation prompt to instruct the LLM to analyze the condition check, as well as agent messages and screenshots from all attempts, in order to produce three aspects of content: (1) where the failure occurred; (2) why the agent failed; (3) what agent behaviors caused the failure. The LLM first refers to the condition check to locate at which step the agent failed. Next, it explains why the agent’s actions were considered failed based on the conditions that were not met. In addition, it summarizes agent messages and webpage screenshots and analyzes what specific actions led to the failure. Lastly, ReUseIt offers actionable tips to help the user provide guidance to unblock the agent’s progress. It assembles the actionable guidance prompt to guide the LLM to summarize the fallback actions at failed steps and outline suggestions on how to help the agent recover for the user’s reference.
@@ 4.3. Implementation
Web agent framework. Our system is built on the Magentic-UI web agent framework (Mozannar et al., 2025a), which offers a web interface that allows human oversight of web agent actions. We chose this web agent framework since its interface displays agent actions on a live webpage with visual overlays (e.g., a rectangle marking the target UI element, a dot showing cursor clicks) so users can monitor execution in real time and spot issues as they arise. In our implementation, we use Magentic-UI’s web surfer  [^2] https://github.com/microsoft/magentic-ui/tree/main/src/magentic_ui/agents/web_surfer for web browsing, interaction, and navigation, as well as the same tool definitions (e.g., visit URL, web search, scroll down/up, click, type)  [^3] https://github.com/microsoft/magentic-ui/tree/main/src/magentic_ui/tools to define the web surfer’s functionality.
LLMs. ReUseIt has been tested with OpenAI GPT-4o (OpenAI, 2024) for webpage parsing and action reasoning of the web surfer. The reasoned actions are performed on the webpage using Playwright (Microsoft, 2025a). The LLMs for workflow synthesis, condition check on webpage states, and user notification were tested with GPT-4o.
Workflow synthesis time. The execution time of each task depends on the number of steps. We report the mean and standard deviation of per-task execution time for the formative study tasks in Table 2. To synthesize the ReUseIt Workflow, each original and variation task is executed five times, so the workflow synthesis time is about the total number of executions multiplied by the per-task execution time. To optimize the workflow synthesis time, multiple executions can be initiated in parallel. Failed and successful executions from these runs can then be used for workflow synthesis.
Table 2. Execution and workflow synthesis time of user study tasks (mins:secs $\pm$ std).
@@
1 Search the ¡flight type¿ ¡cabin type¿ ticket from ¡departure city¿ to ¡destination city¿ on ¡date¿ for ¡number of passengers¿ on ¡website¿ $2{:}38\pm 0{:}41$ $\sim 52{:}40$
2 Go to ¡website¿ to find ¡home filtering¿ ¡home type¿ for ¡rent or buy¿ in ¡city¿. Scrape the information from top ¡num listings¿ $1{:}20\pm 0{:}47$ $\sim 26{:}40$
3 Go to ¡website¿ to find the first ¡num pets¿ ¡pet filtering¿ ¡pet type¿’s ¡pet information¿ that can be adopted near ¡city¿ $1{:}33\pm 0{:}31$ $\sim 31{:}00$
4 Search the lowest price of ¡product¿ with ¡attributes¿ by applying ¡discount¿ on ¡website¿ $1{:}11\pm 0{:}38$ $\sim 23{:}40$
5 Search the ¡paper title¿ on Google Scholar. Find the papers authored by ¡author of interest¿ of the searched article. Retrieve the most recent paper according to ¡sorting criteria¿ and summarize its abstract $0{:}46\pm 0{:}20$ $\sim 15{:}20$
6 Search ¡news topic¿ from ¡website¿, and summarize its author’s recent articles $1{:}09\pm 0{:}39$ $\sim 23{:}00$
This table summarizes the user study tasks, and the time taken to execute the tasks and synthesize workflows.
@@ 5.1. Evaluation Setup
We randomly sampled fifteen tasks from the Skyvern web benchmark (Singh, 2025) for the benchmark evaluation (task details in Table 3). For each task, we automatically generated variation tasks (using the method introduced in Section 4.1), executed the original and variation tasks (a task family) five times, and then synthesized a workflow for each task family. All original tasks had three variations (attribute, category, website) except for Archive (2 variations) and Cars (1 variation), where the missing variations failed in all executions due to captcha errors.
We compared the agent guided by Task + our synthesized ReUseIt Workflow (e.g., Figure 7) against three baselines: Task-Only, Task + Success-Traces, and Task + Magentic-UI Plan. We used the average success rate ($SR$) as the evaluation metric (as introduced in Section 3.2). Following WebVoyager (He et al., 2024), we employed LLM-as-a-judge to assess execution success. Specifically, a GPT-4o evaluator takes the task description, the final three screenshots prior to task termination, and the agent’s answer, and returns a binary success decision accompanied by a rationale. To validate the performance of this LLM judge, we randomly audited 45 judgments (15%) with a human rater. The LLM-judge achieved $\textrm{Accuracy}=0.778$, $\textrm{Precision}=0.852$, $\textrm{Recall}=0.793$, $\textrm{F1}=0.821$. Agreement between the LLM judge and human rater was moderate-substantial with $\textrm{Cohen's }\kappa=0.528$.
@@
This figure shows an example of ”ReUseIt Workflow (User)”. The workflow contains the major actions to take for each step, and for each action, it includes the synthesized condition checks and fallback actions. When user feedback is collected, it is added to the workflow to improve the condition checks or fallback actions.
@@ 5.2. Evaluation Results
Figure 8 presents the benchmark evaluation results for all fifteen task families. We observe that our synthesized workflow has improved the $SR$ by a large margin compared to the baseline methods. Averaging across tasks, the $SR$s are: Task + ReUseIt Workflow ($70.1\%\pm 16.4\%$), Task + Magentic-UI Plan ($48.6\%\pm 12.9\%$), Task + Success-Traces ($41.4\%\pm 14.8\%$), Task-Only ($24.2\%\pm 13.2\%$). ReUseIt Workflow improves $SR$s by about $21.5\%$ over Task + Magentic-UI Plan, $28.7\%$ over Task + Success-Traces, and $45.9\%$ over Task-Only. Appendix D reports additional results on the mean $SR$s with 95% confidence intervals for each and across all fifteen task families.
Figure 8. Benchmark evaluation results of fifteen web automation tasks and their variations (avg $\pm$ std success rate).
This figure shows 15 bar plots with task success rates for the benchmark evaluation tasks.
From the results, we find that providing the agent with successful traces has been effective in improving the $SR$s on the original task, but less effective on the variation tasks compared to Magentic-UI Plan. The largest improvements come from adding execution guards in the ReUseIt Workflow. The condition checks and fallback actions have improved agent reusability by guiding the agent to reflect on actions and recover from failures. These benefits also boosted reliability in executing the same and similar tasks. Additionally, our evaluation spans multiple variation tasks, so higher success rates indicate better generalizability to handle value adjustments and webpage UI differences.
5.3. Ablation Study
To further evaluate the effectiveness of components in our synthesized workflows, we conduct an ablation study to evaluate the effectiveness of fallback actions. In this ablation, we implement a variation named Task + ReUseIt Workflow w/o Fallback Actions: when condition checks are not met, instead of relying on synthesized fallback actions, the web agent needs to explore recovery strategies on its own and retry the failed step (up to three times). We summarize the ablation results of each and across all benchmark evaluation tasks in Table 4. For most tasks, synthesized workflows with fallback actions achieve higher $SR$s than those without. While simple retries give web agents more attempts after failure, unguided retries are less effective than retries guided by the synthesized fallback actions, which enable agents to recover from failures more reliably. Upon close inspection, we noticed that agents were likely to repeat errors they made in their first attempts when trying to fix errors, which highlights that both condition checks and fallback actions are necessary for improving agents’ reliability.
Table 4. Ablation study results. Mean success rates (%) $\uparrow$ $\pm$ std of each task and across tasks.
This table presents the ablation study results. There are three columns in the table, from left to right, the ”Task”, the results of ”Task + ReUseIt Workflow w/o Fallback Actions”, and the results of ”Task + ReUseIt Workflow”.
Task Task + ReUseIt Workflow w/o Fallback Actions Task + ReUseIt Workflow
Archive $46.7\pm 11.6$ $\textbf{86.7}\pm 11.6$
Asus $60.0\pm 16.3$ $\textbf{90.0}\pm 11.6$
Cars $70.0\pm 14.1$ $\textbf{80.0}\pm 0.0$
Bandcamp $60.0\pm 0.0$ $\textbf{90.0}\pm 14.1$
BBC $45.0\pm 19.2$ $\textbf{80.0}\pm 16.3$
Fda $35.0\pm 10.0$ $\textbf{55.0}\pm 10.0$
Forbes $\textbf{45.0}\pm 10.0$ $\textbf{45.0}\pm 30.0$
Getty images $45.0\pm 19.2$ $\textbf{55.0}\pm 10.0$
Groupon $55.0\pm 10.0$ $\textbf{80.0}\pm 16.3$
Restaurantguru $40.0\pm 28.3$ $\textbf{50.0}\pm 41.6$
Scribd $40.0\pm 28.3$ $\textbf{55.0}\pm 19.2$
Smithsonianmag $65.0\pm 19.2$ $\textbf{90.0}\pm 20.0$
Sportskeeda $40.0\pm 28.3$ $\textbf{55.0}\pm 25.2$
Ticketmaster $50.0\pm 20.0$ $\textbf{65.0}\pm 10.0$
Tvguide $55.0\pm 25.2$ $\textbf{75.0}\pm 30.0$
Across Tasks $50.1\pm 10.3$ $\textbf{70.1}\pm 16.4$
@@ 6. User Study
To understand how our system affects the user experience when interacting with web agents, we conducted a user study to explore the following research questions:
@@ 6.1. Study Design
Procedure. The onboarding session ($\sim$10 mins), involving consent, introduction, and an example task walkthrough, followed the same procedure as the formative study. After it, each participant examined two task families covering six agent executions ($\sim$30 mins). We shuffled the task order and ensured each task family was examined by three different participants. In each task, they were provided with a slide deck containing the task description, the screen recording of agent actions, the user notification message, and the workflow showing a sequence of actions the agent followed. After examining all the information, participants were asked to verbally analyze why and how the agent failed or successfully completed the tasks, evaluate the helpfulness of the user notification message to understand the agent’s issues, and (optionally) offer guidance by either writing a short instruction or typing their edits in the workflow. We recorded the desktop screen and transcribed the audio. After the task, participants completed a post-study questionnaire and had a semi-structured interview ($\sim$20 mins). The interview questions can be found in Appendix E.1. Each participant was paid a \$30 gift card to acknowledge their participation.
@@ 6.2. Quantitative Measures
Figure 9. User study follow-up experiment results of six web automation tasks and their variations (avg $\pm$ std success rate). For product price tracking and publication search, users were satisfied with the agent performance in the ReUseIt Workflow condition and did not edit the workflow or offer guidance (i.e., no result of Task + ReUseIt Workflow User is shown).
This figure shows six bar charts with task success rates for the user study follow-up experiment results.
Follow-Up Experiment (S2-RQ1). Figure 9 shows the follow-up experiment results and Appendix E.2.1 presents additional results of mean success rates with 95% confidence intervals for each task and across all tasks. In tasks “Product Price Tracking” and “Publication Search,” participants were satisfied with the ReUseIt Workflow and did not improve it, and we plotted the success rate averaged across only the users who offered guidance in Figure 9 (Task + ReUseIt Workflow User). See Figure 7 for an example workflow updated with user guidance (blue text). Appendix E.2.2 reports the full results for each user. Across tasks, the success rates for Task + ReUseIt Workflow and Task + ReUseIt Workflow User were $86.5\%\pm 9.9\%$ and $87.0\%\pm 7.1\%$. These results show that (1) ReUseIt Workflow can boost the agent’s task execution reliability compared to the baseline methods and (2) user guidance can further help improve the workflow for better agent execution performance on repetitive tasks.
@@
Accompanying the question on required user guidance, we conducted an analysis of user guidance statistics to compare the number and length of user guidance. We received 9, 17, and 17 pieces of user guidance to ReUseIt Workflow, Magentic-UI Plan, and Task-Only, respectively. The number of words (mean $\pm$ std) in user guidance was $43.41\pm 50.02$ (ReUseIt Workflow), $43.24\pm 48.18$ (Magentic-UI Plan), and $56.11\pm 44.59$ (Task-Only). These results indicate that the execution guards in the ReUseIt Workflow have effectively improved the interpretability of agent actions/issues and reduced the guidance burden on users.
Figure 10. User study post-study questionnaire results. These questions collect user feedback on the agent messages’ usefulness to help users understand agent issues and offer guidance, as well as the user-perceived effort on debugging the agent.
This figure presents the responses to the Likert-scale questions from the user study participants. The questions are regarding the agent message’s helpfulness to help users understand agent issues and offer guidance, as well as the user-perceived effort on debugging the agent.
@@ 6.3. Qualitative Findings
From the analysis of interview data, we obtain the following findings centered around S2-RQ2 to S2-RQ4. Overall, we found that the ReUseIt ’s synthesized workflows have helped users to spot and fix agent issues and augmented their willingness to adopt web agents.
Condition checks helped users to identify issues, issue analysis helped them to understand challenges (S2-RQ2). Participants emphasized that ReUseIt Workflow’s condition checks helped them anticipate and localize failures without exhaustive monitoring. S2-P3 noted, “If you know where the model’s gonna fail from the condition checks [in ReUseIt Workflow], you don’t have to watch the agent very carefully and examine everything. You kind of know when the error’s gonna be. Otherwise, there’s always some possibility of making an error at any step. So definitely helps you pinpoint where things are gonna go wrong.” They also commented that the user notification guided by ReUseIt Workflow clarified the reason why the agent failed, which made issue diagnosis easier. For instance, S2-P4 said that “With ReUseIt Workflow, I would understand a little bit better why,” and S2-P3 emphasized the value of this detail: “For ReUseIt Workflow, I feel like even if it gets it wrong, it gives you so much detail that you can exactly intervene on the spot that you want to.”
@@
User notification with challenge identification and guidance tips allowed precise and low-effort user guidance (S2-RQ3). Participants described the user notification provided with ReUseIt Workflow broadly covered the agent’s challenges that supported targeted edits. As S2-P1 explained that “The messages are like a super set of the challenges faced by the agent. I always find that my instruction is a subset of the suggested ideas. And also, I think it’s important that I match the suggested ideas with the actual workflow of the agent because some of the suggested content is already covered by the workflow, so I don’t have to repeat it. Instead, I need to think about why the existing workflow is not good enough.” In addition, S2-P2 underscored the usefulness of actionable guidance, saying that “I do think that the actionable guidance says something about potentially what are some of the next steps, and after reviewing the workflow myself and then editing the prompts, I think the actionable guidance does have a good contribution.” The ReUseIt Workflow, in turn, supported user understanding of agent challenges for them to provide more effective guidance (S2-P1, S2-P8). For example, S2-P8 said that “These [condition checks and fallback actions] actually give me not only hints, but also much guidance on how I should train an agent, or what my approach should be.”
@@
Reliable and controllable web agents improved user willingness for practical use (S2-RQ4). Participants highlighted that the ReUseIt ’s workflows made agent reuse feasible because the agent behaved more consistently and could be steered through explicit condition checks. As S2-P2 put it, “If I wanna reuse it [the agent] reliably, then I would consider being very extensive with my workflow and being very consistent about it… so that I know how it’s supposed to work.” S2-P4 further commented on the connection between user trust and agent reliability, saying, “I think I have a better chance of getting better results just because it [the agent] has… ways of approaching the task, so I think I would trust it more. Especially if I had a certain level of control in terms of the conditions and the fallbacks, I would use it even more just to make sure that the result that I’m getting is the proper one.”